# Determination of the initial condition for the Balitsky-Kovchegov equation with transformers

### v1

@@ 4 Application to DIS observables
Figure 3: Left: Parity plot comparing the transformer emulator predictions for the DIS reduced cross section $\sigma_{r}/(\sigma_{0}/2)$ (y-axis) against the exact dipole model values (x-axis). The dashed line indicates perfect agreement, with mean and median relative errors of $0.115\%$ and $0.073\%$, respectively. Right: Distribution of the signed relative errors, shown within $\pm 1\%$, demonstrating that the overwhelming majority of predictions lie well below the per-mille level. Together these demonstrate the high accuracy of the DIS surrogate model.
With a robust emulator for the BK-evolved dipole amplitude established, we now proceed to its primary application: the calculation of DIS observables for a global analysis of HERA data. This requires convolving the emulated amplitude with the virtual photon wave functions to compute the reduced cross section, $\sigma_{r}$. While these integrals must be evaluated once to generate the training data for the emulator, repeating them for each of the 10 00010\,000 parameter sets across a large number of kinematic points during a global fit would be computationally prohibitive. The surrogate model circumvents this bottleneck by replacing the repeated numerical evaluations with fast predictions, thereby enabling efficient exploration of the parameter space. To overcome this final bottleneck, we construct a second, dedicated transformer based emulator. This network is trained to map directly from the six inputs $(\log_{10}Q_{s0}^{2},\,\gamma,\,e_{c},\,C,\,\log_{10}x,\,\log_{10}Q^{2})$ to the normalized reduced cross section $\sigma_{r}/(\sigma_{0}/2)$. Its architecture follows closely that used for $N(r,x)$, and it is trained on the $\sim 4\times 10^{6}$ values computed from the dipole amplitudes.
For the DIS cross section emulator, we employ a composite loss function that combines mean squared error (MSE) with a Smooth-$L_{1}$ term cite-key ,
@@
This high-speed surrogate enables a full global fit to the combined HERA neutral current $e^{+}p$ data H1:2015ubc . The negligible evaluation cost of the emulator allows for the efficient scanning of millions of points in the parameter space. The best-fit parameters $\bm{\theta}=\bigl(\sigma_{0}/2,\;Q_{s,0}^{2},\;e_{c},\;\gamma,\;C^{2}\bigr)$ are determined by minimizing the chi-squared function
@@
where $\sigma_{r,i}^{\text{th}}(\bm{\theta})$ is the model prediction and $\delta\sigma_{r,i}^{\text{data}}$ denotes the total experimental uncertainty for point $i$. Statistical uncertainties are estimated using the replica method Ball:2008by , in which independent fits are performed on replica data sets generated by adding Gaussian noise to the experimental measurements. This approach has been widely used in the extraction of transverse-momentum-dependent parton distribution functions Moos:2025sal ; Bacchetta:2024qre ; Echevarria:2020hpy . In this work we use 200 replicas to obtain the uncertainty estimates. The minimization is performed using the Minuit algorithm from the iminuit package iminuit .
@@
In our analysis we perform four distinct fits, corresponding to the two choices of the starting point $x_{0}=0.01$ and $x_{0}=0.05$, and to configurations where the anomalous dimension $\gamma$ is either fixed to $1.0$ or treated as a free parameter with $\gamma=1$ corresponding to the standard MV model. Our motivation for considering $x_{0}=0.05$, in addition to the more standard choice $x_{0}=0.01$, is to assess the validity of the small-$x$ framework for more moderate values of $x$. The best fit parameters and corresponding $\chi^{2}$ per degrees of freedom (dof) for all four fit configurations are summarized in Table 1. The overall fit quality is comparable to the recent Bayesian analysis Casuga:2023dcf . For our fits with $x_{0}=0.01$ we obtain $\chi^{2}/\mathrm{dof}$ values close to unity, in line with the results of Ref. Casuga:2023dcf . When the starting point is shifted to $x_{0}=0.05$, the fit quality deteriorates somewhat, as reflected in the larger $\chi^{2}/\mathrm{dof}$ values. Comparing the parameters, we observe that the fitted anomalous dimension $\gamma$ values remain close to 1.0, though the $x_{0}=0.05$ fit prefers a slightly larger value. Differences also appear in the extracted $\sigma_{0}/2$ and $C^{2}$ parameters, reflecting the sensitivity of the fits to the chosen starting point of the BK evolution. The smaller $Q_{s,0}^{2}$ for $x_{0}=0.05$ is expected, as the saturation scale increases with higher energy (i.e. lower $x$).
@@
As a further cross-check, we also examine how the dipole amplitude depends on the choice of the initial rapidity scale. In Fig. 7 we compare the fitted dipole amplitudes $N(r,x)$ obtained from the $x_{0}=0.01$ and $x_{0}=0.05$ fits, evaluated at the same values of $x$. Specifically, we compare the results at $x=10^{-3}$ and $x=10^{-5}$ for both initial conditions $x_{0}=0.01$ and $x_{0}=0.05$. The left panel (linear scale) shows that the two curves nearly coincide, with only a small residual difference around the transition region $r\!\sim\!1\ \mathrm{GeV}^{-1}$. The right panel (log-log scale) demonstrates that in the small-$r$ region the two amplitudes are essentially identical, with deviations well below the percent level. This indicates that once the rapidity shift is properly taken into account, the BK evolution effectively eliminates the dependence on the arbitrary choice of $x_{0}$, leaving only minor residual differences around the transition region. The remaining small discrepancy is compensated by a larger value of $\sigma_{0}/2$ for the $x_{0}=0.05$ fit, such that the produced cross sections are very close to each other when away from the saturation region $N\approx 1$.
@@ 5 Summary
The framework is readily extensible to several crucial areas. Future investigations will include the incorporation of the impact parameter Mantysaari:2024zxq ; Cepila:2018faq ; Cepila:2020xol ; Cepila:2023pvh ; Cepila:2024qge ; Cepila:2025ujv ; Bendova:2019psy , the implementation of next-to-leading order corrections to both the BK kernel Balitsky:2007feb ; Balitsky:2009xg and the photon wave functions Balitsky:2010ze ; Balitsky:2012bs ; Beuf:2011xd ; Beuf:2016wdz ; Beuf:2017bpd ; Hanninen:2017ddy ; Beuf:2021qqa ; Beuf:2021srj ; Beuf:2022ndu , and the application to diffractive DIS measurements both at LO Kowalski:2008sa ; Munier:2003zb ; Marquet:2007nf ; Lappi:2023frf and NLO Beuf:2022kyp ; Beuf:2024msh ; Kaushik:2025roa . A key advantage of our approach is the decoupling of the computationally intensive evolution from the fitting procedure, which permits the straightforward addition of new parameters or the coupling to other physical processes. This work thus provides a robust and flexible tool, paving the way for the next generation of precision global analyses in small-$x$ physics.
@@ Acknowledgment
We thank Bjoern Schenke for helpful discussions and Noah Moran for collaboration during the early stage of this project. Z.K. and J.P. are supported National Science Foundation under grant No. PHY-2515057. D.Y.S. is supported by the National Science Foundations of China under Grant No. 12275052, No. 12147101. This work is also supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration.

### v2

@@ 4 Application to DIS observables
Figure 3: Left: Parity plot comparing the transformer emulator predictions for the DIS reduced cross section $\sigma_{r}/(\sigma_{0}/2)$ (y-axis) against the exact dipole model values (x-axis). The dashed line indicates perfect agreement, with mean and median relative errors of $0.115\%$ and $0.073\%$, respectively. Right: Distribution of the signed relative errors, shown within $\pm 1\%$, demonstrating that the overwhelming majority of predictions lie well below the per-mille level. Together these demonstrate the high accuracy of the DIS surrogate model. We find that $99.755\%$ of the validation points fall within the $\pm 1\%$ interval.
With a robust emulator for the BK-evolved dipole amplitude established, we now proceed to its primary application: the calculation of DIS observables for a global analysis of HERA data. This requires convolving the emulated amplitude with the virtual photon wave functions to compute the reduced cross section, $\sigma_{r}$. While these integrals are evaluated offline to build the training set, repeatedly performing the multi-dimensional integrations for each theory evaluation in an iterative analysis across the full HERA kinematics would be computationally expensive. The surrogate model circumvents this bottleneck by replacing the repeated numerical evaluations with fast predictions, thereby enabling efficient exploration of the parameter space. To overcome this final bottleneck, we construct a second, dedicated transformer based emulator. This network is trained to map directly from the six inputs $(\ln Q_{s0},\,\gamma,\,e_{c},\,\ln C^{2},\,\ln x,\,\ln Q^{2})$ to the normalized reduced cross section $\sigma_{r}/(\sigma_{0}/2)$. Its architecture follows closely that used for $N(r,x)$, and it is trained on the $\sim 4\times 10^{6}$ values computed from the dipole amplitudes.
For the DIS cross section emulator, we employ a composite loss function that combines mean squared error (MSE) with a Smooth-$L_{1}$ term Terven et al. (2025),
@@
This high-speed surrogate enables a full global fit to the combined HERA neutral current $e^{+}p$ data Abramowicz and others (2015). The negligible evaluation cost of the emulator allows for the efficient scanning of millions of points in the parameter space. The best-fit parameters $\bm{\theta}=\bigl(\sigma_{0}/2,\;Q_{s,0}^{2},\;e_{c},\;\gamma,\;C^{2}\bigr)$ are determined by minimizing the chi-squared function
@@
where $\sigma_{r,i}^{\text{th}}(\bm{\theta})$ is the model prediction and $\delta\sigma_{r,i}^{\text{data}}$ denotes the total experimental uncertainty for point $i$. Statistical uncertainties are estimated using the replica method Ball et al. (2009), in which independent fits are performed on replica data sets generated by adding Gaussian noise to the experimental measurements. This approach has been widely used in the extraction of transverse-momentum-dependent parton distribution functions Moos et al. (2025); Bacchetta et al. (2024); Echevarria et al. (2021). In this work we use 200 replicas to obtain the uncertainty estimates. The minimization is performed using the Minuit algorithm from the iminuit package Dembinski and et al. (2020).
We evaluate the $\chi^{2}$ using a diagonal approximation for the experimental covariance. Specifically, the uncertainty in the denominator of Eq. (13) is taken as a point-by-point total uncertainty obtained by adding in quadrature the statistical uncertainty, the uncorrelated systematic uncertainty, and all quoted systematic/procedural contributions provided with the combined HERA dataset. In the present proof-of-principle LO study we therefore do not introduce nuisance parameters for correlated systematics. We leave a full correlated treatment (covariance or nuisance-parameter form) for future work. After applying the kinematic cuts used in this analysis, the fit includes $N_{\rm data}=403$ points for $x_{0}=0.01$ and $N_{\rm data}=425$ points for $x_{0}=0.05$.
@@
In our analysis we perform four distinct fits, corresponding to the two choices of the starting point $x_{0}=0.01$ and $x_{0}=0.05$, and to configurations where the anomalous dimension $\gamma$ is either fixed to $1.0$ or treated as a free parameter with $\gamma=1$ corresponding to the standard MV model. Our motivation for considering $x_{0}=0.05$, in addition to the more standard choice $x_{0}=0.01$, is to assess the validity of the small-$x$ framework for more moderate values of $x$. The best fit parameters and corresponding $\chi^{2}$ per degrees of freedom (dof) for all four fit configurations are summarized in Table 2. The overall fit quality is comparable to the recent Bayesian analysis Casuga et al. (2024). For our fits with $x_{0}=0.01$ we obtain $\chi^{2}/\mathrm{dof}$ values close to unity, in line with the results of Ref. Casuga et al. (2024). When the starting point is shifted to $x_{0}=0.05$, the fit quality deteriorates somewhat, as reflected in the larger $\chi^{2}/\mathrm{dof}$ values. Comparing the parameters, we observe that the fitted anomalous dimension $\gamma$ values remain close to 1.0, though the $x_{0}=0.05$ fit prefers a slightly larger value. Differences also appear in the extracted $\sigma_{0}/2$ and $C^{2}$ parameters, reflecting the sensitivity of the fits to the chosen starting point of the BK evolution. The smaller $Q_{s,0}^{2}$ for $x_{0}=0.05$ is expected, as the saturation scale increases with higher energy (i.e. lower $x$).
@@
As a further cross-check, we also examine how the dipole amplitude depends on the choice of the initial rapidity scale. In Fig. 7 we compare the fitted dipole amplitudes $N(r,x)$ obtained from the $x_{0}=0.01$ and $x_{0}=0.05$ fits, evaluated at the same values of $x$. Specifically, we compare the results at $x=10^{-3}$ and $x=10^{-5}$ for both initial conditions $x_{0}=0.01$ and $x_{0}=0.05$. The left panel (linear scale) shows that the two curves nearly coincide, with only a small residual difference around the transition region $r\!\sim\!1\ \mathrm{GeV}^{-1}$. The right panel (log-log scale) demonstrates that in the small-$r$ region the two amplitudes show a similar behavior, while in the very small-$r$ region (where $N\ll 1$) the relative difference can reach the $\mathcal{O}(10\%)$ level. This indicates that once the rapidity shift is properly taken into account, the BK evolution effectively eliminates the dependence on the arbitrary choice of $x_{0}$, leaving only minor residual differences around the transition region. The remaining small discrepancy is compensated by a larger value of $\sigma_{0}/2$ for the $x_{0}=0.05$ fit, such that the produced cross sections are very close to each other when away from the saturation region $N\approx 1$.
@@
Finally, we summarize the computational cost of the present workflow to quantify the speed-up enabled by the surrogate strategy. In our current implementation, constructing the training library requires solving the BK equation for $10^{4}$ sampled parameter configurations, with a walltime of about $10$ minutes per configuration on a single CPU core, corresponding to a total of $\sim 1.7\times 10^{3}$ core-hours. Training the DIS cross section transformer on a single NVIDIA A100 GPU takes about $40$ hours. Once trained, the full fit pipeline—including the evaluation of $\sigma_{r}$ over the HERA kinematics and the subsequent minimization—typically completes in about $2$ minutes. These benchmarks demonstrate that the emulator removes the dominant bottlenecks of repeated BK evolutions and multi-dimensional integrations in iterative fits, thereby enabling efficient exploration of the parameter space at negligible per-evaluation cost.
@@ 5 Summary
The framework is readily extensible to several crucial areas. Future investigations will include the incorporation of the impact parameter Mäntysaari et al. (2025); Cepila et al. (2019, 2020, 2024, 2025a, 2025b); Bendova et al. (2019), the implementation of next-to-leading order corrections to both the BK kernel Balitsky and Chirilli (2008, 2009) and the photon wave functions Balitsky and Chirilli (2011, 2013); Beuf (2012, 2016, 2017); Hänninen et al. (2018); Beuf et al. (2021, 2022b, 2022c), and the application to diffractive DIS measurements both at LO Kowalski et al. (2008); Munier and Shoshi (2004); Marquet (2007); Lappi et al. (2023) and NLO Beuf et al. (2022a, 2024); Kaushik et al. (2025). A key advantage of our approach is the decoupling of the computationally intensive evolution from the fitting procedure, which permits the straightforward addition of new parameters or the coupling to other physical processes. This work thus provides a robust and flexible tool, paving the way for the next generation of precision global analyses in small-$x$ physics.
Appendix A Transformer nomenclature used in this work
Our emulator adopts an encoder-style Transformer with self-attention Vaswani et al. (2017), layer normalization Ba et al. (2016), and the pre-normalization (Pre-LN) layout commonly used for stable optimization Xiong et al. (2020). We summarize here the terminology used in Sec. 3 and its concrete meaning in our BK-emulator setup.
In this work, one emulator query (and one training sample) corresponds to one input vector
$\mathbf{x}=\bigl(\ln Q_{s0},\,e_{c},\,\gamma,\,\ln C^{2},\,\ln(r/{\rm GeV}^{-1}),\,\ln(x_{0}/x)\bigr),$ (14)
which specifies one evaluation point of the target function $N(r,x;\bm{\theta})$ with $\bm{\theta}=(Q_{s0},e_{c},\gamma,C^{2})$. The logarithms are used only to improve numerical conditioning; the unit conventions follow those in Sec. 3. Each entry $x_{i}\in\mathbb{R}$ of $\mathbf{x}$ is a scalar feature (a single real number). We map each scalar to a $d_{\rm model}$-dimensional token (embedding vector) using a shared affine map implemented as a linear layer applied feature-wise:
$\mathbf{t}_{i}=W\,x_{i}+\mathbf{b}\,,\qquad i=1,\dots,6,$ (15)
[…]
Because self-attention alone does not encode the ordering of tokens, we add a positional encoding to each token before the first encoder layer, following the standard sinusoidal construction introduced with the Transformer Vaswani et al. (2017). Denoting the positional encoding by $\mathbf{p}_{i}$, the input to the encoder stack is $\mathbf{h}_{i}^{(0)}=\mathbf{t}_{i}+\mathbf{p}_{i}$.
Self-attention updates each token by allowing it to interact with all other tokens in the sequence. In scaled dot-product attention Vaswani et al. (2017), one forms queries, keys and values $Q=HW_{Q}$, $K=HW_{K}$, $V=HW_{V}$ from the matrix $H$ whose rows are the token vectors, and computes
${\rm Attention}(Q,K,V)={\rm softmax}\!\left(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}}\right)V\,.$ (16)
Here $H\in\mathbb{R}^{L\times d_{\rm model}}$ stacks the $L=6$ token vectors row-wise; in multi-head attention the projections are performed separately in each head, with $d_{k}=d_{\rm model}/n_{\rm head}$. Multi-head attention repeats this operation in parallel with multiple learned projections and concatenates the results, enabling the model to capture different types of correlations between input components (here, between parameters and kinematic variables).
A Transformer encoder block consists of a multi-head self-attention sub-layer and a position-wise feed-forward network (FFN), each wrapped with residual connections and layer normalization. We use a Pre-LN layout, i.e. the layer normalization is applied inside each residual branch:
$\tilde{H}$ $=H+{\rm MHA}\bigl({\rm LN}(H)\bigr),$ (17)
[…]
where $\rm LN$} is layer normalization Ba et al. (2016). The role of Pre-LN versus Post-LN in optimization stability is discussed in Ref. Xiong et al. (2020).
After stacking several encoder layers we obtain final token representations $\{\mathbf{h}_{i}\}$. We then apply mean pooling $\bar{\mathbf{h}}=\frac{1}{6}\sum_{i=1}^{6}\mathbf{h}_{i}$ and feed $\bar{\mathbf{h}}$ to an MLP head to produce a single real-valued logit. Here “logit” denotes the real-valued output before the final sigmoid nonlinearity. Applying a sigmoid to the logit yields the dipole amplitude $N(r,x)\in(0,1)$, which automatically enforces the unitarity bound.
@@ Acknowledgment
We thank Bjoern Schenke for helpful discussions and Noah Moran for collaboration during the early stage of this project. M.S.G. and D.Y.S. are supported by the National Science Foundations of China under Grant No. 12275052, No. 12147101, and Shanghai Education Committee under Grant No. 24KXZNB04. Z.K. and J.P. are supported National Science Foundation under grant No. PHY-2515057. This work is also supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics, within the framework of the Saturated Glue (SURGE) Topical Theory Collaboration.